# Lindblad-like quantum tomography for non-Markovian quantum dynamical maps

### v1

@@ III.3.1 Markovian semi-classical dephasing
In order to find the optimal evolution times $t_{1},t_{2}$ that maximize the precision of our estimates, we can minimize the covariance in Eq. (31) which, in turn, requires maximizing the Fisher information matrix (30). By Taylor expanding the cost function, we can actually find a linear relation between the estimate difference $\delta{\bm{\hat{\theta}}}=\bm{\hat{\theta}}_{{\rm F}}-\bm{\theta}_{\star}$, and the differences between the parametrized probabilities and the relative frequencies $\delta\tilde{f}_{i}=p_{i}^{\rm TL}(m_{x}|\bm{\hat{\theta}})-\tilde{f}_{i}(m_{x}\,|\bm{\theta}_{\star})$, namely
@@
where $\mathcal{N}=\Gamma^{\prime}_{T_{2}}\!(t_{1})\Gamma^{\prime}_{\tau_{\rm c}}\!(t_{2})-\Gamma^{\prime}_{\tau_{\rm c}}\!(t_{1})\Gamma^{\prime}_{T_{2}}\!(t_{2})$, and we have introduced a shorthand notation for the partial derivatives $\Gamma^{\prime}_{\tau}\!(t_{i})=\partial\Gamma(t_{i})/\partial\tau_{\rm c}$, $\Gamma^{\prime}_{T_{2}}\!(t_{i})=\partial\Gamma(t_{i})/\partial T_{2}$. The first thing one notices is that, fixing $t_{2}=t_{1}$, the factor $\mathcal{N}=0$, and the difference between the estimation and the true value of the parameters diverges, signaling the fact that one cannot learn two noise parameters using a Ramsey estimator with a single instant of time. The second result one finds is that, in the asymptotic limit $N_{i}\gg 1$, the estimate differences will follow a bi-variate normal distribution. This follows from the fact that Eq. (42) is a linear combination of the differences between the finite frequencies and the binomial probabilities, which are known to follow a normal distribution $N(\bm{0},{\rm diag}\big{(}\sigma^{2}_{f_{1}},\sigma^{2}_{f_{2}})\big{)}$ with binomial variances $\sigma^{2}_{f_{i}}$ defined in Eq. (33). Therefore, the frequentist estimates will also be normally distributed $\delta{\bm{\hat{\theta}}}\sim N(\bm{0},\Sigma_{\bm{\hat{\theta}}}\big{)}$ according to
@@
where $M_{\bm{\hat{\theta}}}(t_{1},t_{2})$ is the matrix in Eq. (42). A more detailed derivation of the relationship between shot noise and the uncertainty in estimation, as well as the asymptotic covariance matrix, is provided in App. D. From this perspective, the aforementioned divergence for $t_{1}=t_{2}$ is a consequence of the singular nature of this matrix, which cannot be thus inverted.
@@
finding the two optimal values at which the signal shows the highest sensitivity to changes in the OU noise (see Fig. 3). The optimal times obtained in this way are depicted in Fig. 4 as a function of the noise correlation time. In the limit where this correlation time is much smaller than the effective decoherence time $\tau_{\rm c\star}\ll T_{2\star}$, the signal only carries important information about the noise for times that are much larger than the correlation time. Hence, we are in the long-time limit where the decay rate is constant $\gamma(t)\approx 1/2T_{2}$ and one expects to find agreement with a purely Lindbladian dephasing noise. As discussed in App. B, the LQT for pure dephasing requires a single measuring time, and can be analytically found by minimizing the standard deviation of the estimated noise parameter (65). This solution yields an optimal time $t_{\rm opt}=0.797T_{2\star}$, which is actually very close to the intercept of the curve of $t_{2,{\rm opt}}$ shown in Fig. 4. In this long-time regime, we find $t_{1,{\rm opt}}\approx 0$ indicating that measurements at time $t_{2,\rm opt}$ will be mainly used to determine parameter $T_{2}$, while those at $t_{1,{\rm opt}}\approx 0$ contribute to determine the much smaller $\tau_{\rm c}$.
@@
Figure 4: Optimal times for OU dephasing $\ell$QT: We represent $(t_{1,\rm opt},t_{2,\rm opt})=\texttt{argmin}\{\det(\Sigma(t_{1},t_{2}))\}$ as a function of the ratio of the real noise parameters $\tau_{\rm c\star}/T_{2\star}$. In the regime where $\tau_{\rm c\star}\ll T_{2\star}$, we find that $t_{2}\approx 0.8T_{2\star}$ and $t_{1}\approx 0$. The optimal time of this measurement lies very close to the Lindbladian result $t=0.797T_{2\star}$ found when there is no time correlation and we have a single parameter $T_{c}$, which is an exact analytical result discussed in App. B.
Bayesian Ramsey estimators.
Let us now describe how a Bayesian inference for OU dephasing $\ell$QT would proceed, which will provide an experimental procedure to operate at the optimal estimation times. We start by commenting on the fully-uncorrelated Lindbladian limit discussed in App. B, where the optimization of the measurement time for each Bayesian step (37) can also be solved analytically. Considering that the prior probability distribution $\pi_{\ell}(\gamma)$ for our knowledge about the decay rate at the $\ell$-th step is Gaussian, we can focus on how its mean and variance change as one takes the next Bayesian step. In the Appendix, we show that, minimizing the Bayesian variance (68) of the next step, one finds optimal measurement times that agree with the above frequentist prediction, albeit for the knowledge of the decay rate that we actually have at each particular step $t_{\ell}=0.797/2\hat{\gamma}_{\ell}$ or, alternatively, of the decoherence time $T_{2,\ell}$. This result is very encouraging, as the experimentalist may only have a crude guess of this value, but it gets automatically updated towards the optimal regime. This motivates an extension to time-correlated dephasing such as the OU noise.
@@
For the OU dephasing, we have two parameters to learn, and we can maximize the Kullback-Leibler divergence (37) to obtain the subsequent optimal time $t_{\ell}$ for the next Ramsey measurement(s), and the corresponding extension of the data set $\mathbb{D}_{\ell-1}\mapsto\delta\mathbb{D}_{\ell}=\mathbb{D}_{\ell-1}\cup\delta\mathbb{D}_{\ell-1}$. We then proceed by measuring at this time, updating the prior, and starting the optimization step all over again to finally find the estimates (35) $\bm{\hat{\theta}}_{{\rm B}}=(\hat{\tau}_{{\rm c},\ell},\hat{T}_{2,\ell})$. As shown in Fig. 5, as one collects more and more data, the Bayesian measurement times cluster at two single times, and tend to alternate between them. Remarkably, these times are the optimal $t_{1,{\rm opt}}$ and $t_{2,{\rm opt}}$ predictions of the frequentist approach shown in Fig. 4. We can see how the Bayesian approach automatically finds the optimal measurement setting to learn a time-correlated dephasing noise.
@@
III.3.2 Non-Markovian quantum dephasing
@@
We consider an apparently mild modification of the noise PSD with respect to the OU case (39). In particular, we use
@@
which is a Lorentzian of width $\kappa$ centered around $-\Delta_{\rm c}$, and reaching a maximum of $16g_{\bar{n}}^{2}/\kappa$. We note that for $\Delta_{\rm c}=0$, we recover the previous OU case (39) with $\tau_{\rm c}=2/\kappa$ and $c=4g^{2}_{\bar{n}}\kappa$. On the other hand, for $\Delta_{\rm c}\neq 0$, this PSD is not an even function $S(\omega)\neq S(-\omega)$, and the associated frequency noise cannot arise from a semi-classical stochastic model [113]. Instead, this particular PSD can be deduced from a quantum-mechanical dephasing model as discussed in App. A, and applied to a qubit coupled to a dissipative bosonic mode.
@@
In any of the two architectures discussed, when the coupling between the bosonic mode and the qubit is weaker than the dissipative rate $g_{\bar{n}}\ll\kappa$, one can truncate the cumulant expansion of a time-convolutionless master equation of the qubit at second order such that, after tracing over the driven-dissipative mode in its stationary state, one arrives at a time-local dephasing master equation of the form (12). This master equation will be controlled by an auto-correlation function (63) for the bath operator $B(t)=2g(a^{\dagger}a-\bar{n})$, following the notation used below Eq. (13) and in App. A. In particular, making use of the quantum regression theorem [119], this auto-correlation can be expressed as
@@
which coincides with the OU auto-correlation function (38) when $\Delta_{\rm c}=0$. Being wide-sense stationary, one can Fourier transform this function (14), leading to the displaced Lorentzian PSD in Eq. (46). Following Eq. (15), one can obtain the following time-dependent decay rate
@@
Figure 7: Non-Markovianity measure for the quantum dephasing map: non-Markovianity measure of Eq. (24) for different values of $\tau_{c}=2/\kappa$ and $\Delta_{\rm c}$. The dashed line signals the limit between the Markovian and the non-Markovian regions and corresponds to the equation $\Delta_{\rm c}\approx 1.82\kappa$.
@@
Frequentist Ramsey estimators.
As in the OU case, we need to minimize $\mathsf{C}^{\rm pd}_{\mathrm{TL}}(\bm{\theta})$ in Eq. (26), where the likelihood function $p^{\rm TL}_{i}(m_{x}|\bm{\theta})=\big{(}1+(-1)^{m_{x}}\mathord{\rm e}^{-\Gamma(t_{i})}\big{)}/2$ now depends on the new attenuation factor (49). Since there are three noise parameters $\bm{\theta}_{\star}=(g^{2}_{\bar{n}\star},\kappa_{\star},\Delta_{{\rm c}\star})\in\Theta=\mathbb{R}^{3}_{+}$, we shall at least need to measure at three different times. To assess the performance of the frequentist approach under shot noise, we numerically generate the relative frequencies $\tilde{f}_{1}(m_{x}\,|\bm{\theta}_{\star}),\tilde{f}_{2}(m_{x}\,|\bm{\theta}_{\star}),\tilde{f}_{3}(m_{x}\,|\bm{\theta}_{\star})$ at these times by sampling the probability distribution with the real noise parameters $\bm{\theta}_{\star}$ a number of times $N_{\rm shot}=N_{1}+N_{2}+N_{3}$. In order to find the three optimal times, we minimize the determinant of the covariance matrix (44), which will have an similar expression as Eq. (43), but now expressed in terms of $3\times 3$ matrices for the underlying trivariate normal distribution. Since we are optimizing $t_{1}$, $t_{2}$ and $t_{3}$, one may wonder if we also could improve the estimation by redistributing the total number of measurements differently at each of these times. However, according to our prescription in which the imprecision is quantified by the long-run Gaussian pdf, which defines an elliptical volume in this case, one gets $\det\Sigma_{\bm{\hat{\theta}}}(t_{1,\rm opt},t_{2,\rm opt},t_{3,\rm opt})\propto{1}/{N_{1}N_{2}N_{3}}$. Therefore, the number of measurements must be equally distributed between the three different times $N_{1}=N_{2}=N_{3}=N_{\rm shot}/3$.
@@
Bayesian Ramsey estimators.
Let us now move to the Bayesian approach, where we have some prior knowledge (34) of the parameters $\bm{\theta}=(g^{2}_{\bar{n}},\kappa,\Delta_{\rm c})$ that gets updated at each Bayesian step $t_{\ell}$ by enlarging the data set with $\delta\mathbb{D}_{\ell}$. Minimizing the relative entropy between the prior and the posterior (37), we obtain the subsequent evolution time $t_{\ell}$, and update our knowledge about the noise parameters in the best possible way. As shown in Fig. 9, the Bayesian procedure starts by using evolution update times that are scattered in a broad range of values. However, as the number of iterations increases and we gain more knowledge, they tend to cluster around three well-defined times. In fact, as shown by the corresponding dashed lines, these times coincide with the optimal measurement times of the frequentist approach in Fig. 8. Therefore, if we start a Bayesian experiment with some prior, and we let the experiment run for sufficiently long number of steps, we will learn the optimal times automatically.
@@
In order to compare the precision of the frequentist and Bayesian approaches, we proceed in analogy to the OU noise by looking for a parameter that captures the relative precision of the two approaches (45). We now have to consider that the determinant of the trivariate Gaussian covariance $\det\Sigma$ is a volume, and we can define an average radius as $\bar{R}=\sqrt[3]{\det\Sigma}/(4\pi/3)$. We can quantify the precision by taking the square root of this radius, which scales like a standard deviation. Altogether, the relative precision of the frequentist and Bayesian approaches is defined by the ratio
@@
Figure 10: Ratio between determinants of covariance matrices of frequentist and Bayesian approaches: The ratio of Eq. (50) computed for different parameter values $\tau_{\mathrm{c}\star}$ and $\Delta_{\mathrm{c}\star}$, and a fixed number of total measurements $N_{\rm shot}=2\cdot 10^{4}$. The dashed white line separates the Markovian from the non-Markovian regime. The continuous white line indicates the level where the ratio equals 1. 5000 frequentist runs were done to estimate $\det\mathrm{Cov(\hat{\bm{\theta}}_{\mathrm{F,opt}})}$, while 30 Bayesian runs were done to estimate $\det\Sigma_{\ell_{\rm max}}$ for each point $(\tau_{\rm c\star},\Delta_{\rm c\star})$. In the Bayesian approach 100 shots were taken at each step and the prior is a continuous uniform distribution with $\bm{\theta}\in[\bm{\theta}_{\star}/3,3\bm{\theta}_{\star}]$. Analogously, for the least-squares minimization used in the frequentist approach (trust-region reflective algorithm), we set the same parameter bounds as the ones of the uniform distribution.
@@
It is also useful to quantify the advantage of one estimator with respect to the other in terms of number of measurements that is required to reach a target precision, which amounts to reaching the same value of the above covariance determinant. The values shown for the ratio of the determinants in Fig. 10 can be converted into the ratio of number of measurements by assuming that both covariance determinants in Eq. (50) scale as $1/\sqrt{N_{\rm shot}}$ even in the non-asymptotic regime. Although there can be corrections to this scaling, this can give us in most cases an idea of the proportion of measurement shots one can save by using the best estimator. With this assumption, we get $N_{\rm shot,F}/N_{\rm shot,B}\approx r_{\rm NM}^{2}$, with $r_{\rm NM}$ the ratio of the determinants of two different approaches. Thus, if for instance $r_{\mathrm{\rm NM}}=1.5$, we obtain that $N_{\rm shot,F}/N_{\rm short,B}\approx 2.25$. Therefore, we get more than a 2×2$\times$2 × reduction in the number of measurements with the Bayesian approach in comparison to the frequentist one to reach the same precision. Going back to the values of Fig. 10, we see that the Bayesian approach can result in a considerable improvement as one goes deep in the non-Markovian limit. Before closing this subsection, it is worth recalling once more that, in practice, the frequentist estimation will never be performed at the three optimal times, and the advantage of the Bayesian approach can even be bigger. In App. C, we present a detailed comparison of these estimators with another one in which the shots are evenly distributed between the measurements after evolution times that cover uniformly the whole time interval $T$. As discussed in the appendix, the Bayesian approach is preferable for most of the parameter values, and can again show a big advantage as one enters into the non-Markovian regime.
IV Conclusions and outlook

### v2

@@ Results
In order to find the optimal evolution times $t_{1},t_{2}$ that maximize the precision of our estimates, we can minimize the covariance in Eq. (13) which, in turn, requires maximizing the Fisher information matrix in Eq. (12). By Taylor expanding the cost function, we can actually find a linear relation between the estimate difference $\delta{\bm{\hat{\theta}}}=\bm{\hat{\theta}}_{{\rm F}}-\bm{\theta}_{\star}$, and the differences between the parametrized probabilities and the relative frequencies $\delta\tilde{f}_{i}=p_{i}^{\rm TL}(m_{x}|\bm{\hat{\theta}})-\tilde{f}_{i}(m_{x}\,|\bm{\theta}_{\star})$, namely
@@
where $\mathcal{N}=\Gamma^{\prime}_{T_{2}}\!(t_{1})\Gamma^{\prime}_{\tau_{\rm c}}\!(t_{2})-\Gamma^{\prime}_{\tau_{\rm c}}\!(t_{1})\Gamma^{\prime}_{T_{2}}\!(t_{2})$, and we have introduced a shorthand notation for the partial derivatives $\Gamma^{\prime}_{\tau}\!(t_{i})=\partial\Gamma(t_{i})/\partial\tau_{\rm c}$, $\Gamma^{\prime}_{T_{2}}\!(t_{i})=\partial\Gamma(t_{i})/\partial T_{2}$. The first thing one notices is that, fixing $t_{2}=t_{1}$, the factor $\mathcal{N}=0$, and the difference between the estimation and the true value of the parameters diverges, signaling the fact that one cannot learn two noise parameters using a Ramsey estimator with a single instant of time. The second result one finds is that, in the asymptotic limit $N_{i}\rightarrow\infty$, the estimate differences will follow a bi-variate normal distribution. This follows from the fact that Eq. (39) is a linear combination of the differences between the finite frequencies and the binomial probabilities, which are known to follow a normal distribution $N(\bm{0},{\rm diag}\big{(}\sigma^{2}_{f_{1}},\sigma^{2}_{f_{2}})\big{)}$ with binomial variances $\sigma^{2}_{f_{i}}$ defined in Eq. (15). Therefore, the frequentist estimates will also be normally distributed $\delta{\bm{\hat{\theta}}}\sim N(\bm{0},\Sigma_{\bm{\hat{\theta}}}\big{)}$ according to
@@
where $M_{\bm{\hat{\theta}}}(t_{1},t_{2})$ is the matrix in Eq. (39). A more detailed derivation of the relationship between shot noise and the uncertainty in estimation, as well as the asymptotic covariance matrix, is provided in the Methods section. From this perspective, the aforementioned divergence for $t_{1}=t_{2}$ is a consequence of the singular nature of this matrix, which cannot be thus inverted.
@@
finding the two optimal values at which the signal shows the highest sensitivity to changes in the OU noise (see Fig. 3). The optimal times obtained in this way are depicted in Fig. 4 as a function of the noise correlation time. In the limit where this correlation time is much smaller than the effective decoherence time $\tau_{\rm c\star}\ll T_{2\star}$, the signal only carries important information about the noise for times that are much larger than the correlation time. Hence, we are in the long-time limit where the decay rate is constant $\gamma(t)\approx 1/2T_{2}$ and one expects to find agreement with a purely Lindbladian dephasing noise. As discussed in Sec. I of the Supplementary Material, the LQT for pure dephasing requires a single measuring time, and can be analytically found by minimizing the standard deviation of the estimated noise parameter. This solution yields an optimal time $t_{\rm opt}=0.797T_{2\star}$, which is actually very close to the intercept of the curve of $t_{2,{\rm opt}}$ shown in Fig. 4. In this long-time regime, we find $t_{1,{\rm opt}}\approx 0$ indicating that measurements at time $t_{2,\rm opt}$ will be mainly used to determine parameter $T_{2}$, while those at $t_{1,{\rm opt}}\approx 0$ contribute to determine the much smaller $\tau_{\rm c}$.
@@
Figure 4: Optimal times for OU dephasing $\ell$QT: We represent $(t_{1,\rm opt},t_{2,\rm opt})=\texttt{argmin}\{\det(\Sigma(t_{1},t_{2}))\}$ as a function of the ratio of the real noise parameters $\tau_{\rm c\star}/T_{2\star}$. In the regime where $\tau_{\rm c\star}\ll T_{2\star}$, we find that $t_{2}\approx 0.8T_{2\star}$ and $t_{1}\approx 0$. The optimal time of this measurement lies very close to the Lindbladian result $t=0.797T_{2\star}$ (gray dashed line) found when there is no time correlation and we have a single parameter $T_{c}$, which is an exact analytical result discussed in Sec. I of the Supplementary Material.
b. Bayesian Ramsey estimators. Let us now describe how a Bayesian inference for OU dephasing $\ell$QT would proceed, which will provide an experimental procedure to operate at the optimal estimation times. We start by commenting on the fully-uncorrelated Lindbladian limit discussed in Sec. I of the Supplementary Material, where the optimization of the measurement time for each Bayesian step in Eq. (19) can also be solved analytically. Considering that the prior probability distribution $\pi_{\ell}(\gamma)$ for our knowledge about the decay rate at the $\ell$-th step is Gaussian, we can focus on how its mean and variance change as one takes the next Bayesian step. In Sec. I of the Supplementary Material, we show that, minimizing the Bayesian variance of the next step, one finds optimal measurement times that agree with the above frequentist prediction, albeit for the knowledge of the decay rate that we actually have at each particular step $t_{\ell}=0.797/2\hat{\gamma}_{\ell}$ or, alternatively, of the decoherence time $T_{2,\ell}$. This result is very encouraging, as the experimentalist may only have a crude guess of this value, but it gets automatically updated towards the optimal regime. This motivates an extension to time-correlated dephasing such as the OU noise.
@@
For the OU dephasing, we have two parameters to learn, and we can maximize the Kullback-Leibler divergence of Eq. (19) to obtain the subsequent optimal time $t_{\ell}$ for the next Ramsey measurement(s), and the corresponding extension of the data set $\mathbb{D}_{\ell-1}\mapsto\delta\mathbb{D}_{\ell}=\mathbb{D}_{\ell-1}\cup\delta\mathbb{D}_{\ell-1}$. We then proceed by measuring at this time, updating the prior, and starting the optimization step all over again to finally find the estimates in Eq. (17) $\bm{\hat{\theta}}_{{\rm B}}=(\hat{\tau}_{{\rm c},\ell},\hat{T}_{2,\ell})$. As shown in Fig. 5, as one collects more and more data, the Bayesian measurement times cluster at two single times, and tend to alternate between them. Remarkably, these times are the optimal $t_{1,{\rm opt}}$ and $t_{2,{\rm opt}}$ predictions of the frequentist approach shown in Fig. 4. We can see how the Bayesian approach automatically finds the optimal measurement setting to learn a time-correlated dephasing noise.
@@
Non-Markovian quantum dephasing
@@
We consider an apparently mild modification of the noise PSD with respect to the OU case in Eq. (36). In particular, we use
@@
which is a Lorentzian of width $\kappa$ centered around $-\Delta_{\rm c}$, and reaching a maximum of $16g_{\bar{n}}^{2}/\kappa$. We note that for $\Delta_{\rm c}=0$, we recover the previous OU case (36) with $\tau_{\rm c}=2/\kappa$ and $c=4g^{2}_{\bar{n}}\kappa$. On the other hand, for $\Delta_{\rm c}\neq 0$, this PSD is not an even function $S(\omega)\neq S(-\omega)$, and the associated frequency noise cannot arise from a semi-classical stochastic model [123]. Instead, this particular PSD can be deduced from a quantum-mechanical dephasing model as discussed in the Methods section, and applied to a qubit coupled to a dissipative bosonic mode.
@@
In any of the two architectures discussed, when the coupling between the bosonic mode and the qubit is weaker than the dissipative rate $g_{\bar{n}}\ll\kappa$, one can truncate the cumulant expansion of a time-convolutionless master equation of the qubit at second order such that, after tracing over the driven-dissipative mode in its stationary state, one arrives at a time-local dephasing master equation of the form given in Eq. (20). This master equation will be controlled by an auto-correlation function for the bath operator $B(t)=2g(a^{\dagger}a-\bar{n})$, following the notation used below Eq. (21) and in the Methods section. In particular, making use of the quantum regression theorem [129], this auto-correlation can be expressed as
@@
which coincides with the OU auto-correlation function in Eq. (35) when $\Delta_{\rm c}=0$. Being wide-sense stationary, one can Fourier transform this function as shown in Eq. (22), leading to the displaced Lorentzian PSD in Eq. (43). Following Eq. (23), one can obtain the following time-dependent decay rate
@@
Figure 7: Non-Markovianity measure for the quantum dephasing map: non-Markovianity measure of Eq. (32) for different values of $\tau_{c}=2/\kappa$ and $\Delta_{\rm c}$. The dashed line signals the limit between the Markovian and the non-Markovian regions and corresponds to the equation $\Delta_{\rm c}\approx 1.82\kappa$. The non-Markovian region is located in the upper right corner.
@@
a. Frequentist Ramsey estimators. As in the OU case, we need to minimize $\mathsf{C}^{\rm pd}_{\mathrm{TL}}(\bm{\theta})$ in Eq. (34), where the likelihood function $p^{\rm TL}_{i}(m_{x}|\bm{\theta})=\big{(}1+(-1)^{m_{x}}\mathord{\rm e}^{-\Gamma(t_{i})}\big{)}/2$ now depends on the new attenuation factor in Eq. (46). Since there are three noise parameters $\bm{\theta}_{\star}=(g^{2}_{\bar{n}\star},\kappa_{\star},\Delta_{{\rm c}\star})\in\Theta=\mathbb{R}^{3}_{+}$, we shall at least need to measure at three different times. To assess the performance of the frequentist approach under shot noise, we numerically generate the relative frequencies $\tilde{f}_{1}(m_{x}\,|\bm{\theta}_{\star}),\tilde{f}_{2}(m_{x}\,|\bm{\theta}_{\star}),\tilde{f}_{3}(m_{x}\,|\bm{\theta}_{\star})$ at these times by sampling the probability distribution with the real noise parameters $\bm{\theta}_{\star}$ a number of times $N_{\rm shot}=N_{1}+N_{2}+N_{3}$. In order to find the three optimal times, we minimize the determinant of the covariance matrix in Eq. (41), which will have an similar expression as Eq. (40), but now expressed in terms of $3\times 3$ matrices for the underlying trivariate normal distribution. Since we are optimizing $t_{1}$, $t_{2}$ and $t_{3}$, one may wonder if we also could improve the estimation by redistributing the total number of measurements differently at each of these times. However, according to our prescription in which the imprecision is quantified by the long-run Gaussian PDF, which defines an elliptical volume in this case, one gets $\det\Sigma_{\bm{\hat{\theta}}}(t_{1,\rm opt},t_{2,\rm opt},t_{3,\rm opt})\propto{1}/{N_{1}N_{2}N_{3}}$. Therefore, the number of measurements must be equally distributed between the three different times $N_{1}=N_{2}=N_{3}=N_{\rm shot}/3$.
@@
b. Bayesian Ramsey estimators. Let us now move to the Bayesian approach, where we have some prior knowledge as shown in Eq. (16) of the parameters $\bm{\theta}=(g^{2}_{\bar{n}},\kappa,\Delta_{\rm c})$ that gets updated at each Bayesian step $t_{\ell}$ by enlarging the data set with $\delta\mathbb{D}_{\ell}$. Minimizing the relative entropy between the prior and the posterior in Eq. (19), we obtain the subsequent evolution time $t_{\ell}$, and update our knowledge about the noise parameters in the best possible way. As shown in Fig. 9, the Bayesian procedure starts by using evolution update times that are scattered in a broad range of values. However, as the number of iterations increases and we gain more knowledge, they tend to cluster around three well-defined times. In fact, as shown by the corresponding dashed lines, these times coincide with the optimal measurement times of the frequentist approach in Fig. 8. Therefore, if we start a Bayesian experiment with some prior, and we let the experiment run for sufficiently long number of steps, we will learn the optimal times automatically.
@@
In order to compare the precision of the frequentist and Bayesian approaches, we proceed in analogy to the OU noise by looking for a parameter that captures the relative precision of the two approaches in Eq. (42). We now have to consider that the determinant of the trivariate Gaussian covariance $\det\Sigma$ is a volume, and we can define an average radius as $\bar{R}=\sqrt[3]{\det\Sigma}/(4\pi/3)$. We can quantify the precision by taking the square root of this radius, which scales like a standard deviation. Altogether, the relative precision of the frequentist and Bayesian approaches is defined by the ratio
@@
Figure 10: Ratio between determinants of covariance matrices of frequentist and Bayesian approaches: The ratio of Eq. (47) computed for different parameter values $\tau_{\mathrm{c}\star}$ and $\Delta_{\mathrm{c}\star}$, and a fixed number of total measurements $N_{\rm shot}=2\cdot 10^{4}$. The dashed white line separates the Markovian from the non-Markovian regime, with the non-Markovian region located in the upper right corner. The continuous white line indicates the level where the ratio equals 1. 5000 frequentist runs were done to estimate $\det\mathrm{Cov(\hat{\bm{\theta}}_{\mathrm{F,opt}})}$, while 30 Bayesian runs were done to estimate $\det\Sigma_{\ell_{\rm max}}$ for each point $(\tau_{\rm c\star},\Delta_{\rm c\star})$. In the Bayesian approach 100 shots were taken at each step and the prior is a continuous uniform distribution with $\bm{\theta}\in[\bm{\theta}_{\star}/3,3\bm{\theta}_{\star}]$. Analogously, for the least-squares minimization used in the frequentist approach (trust-region reflective algorithm), we set the same parameter bounds as the ones of the uniform distribution.
@@
It is also useful to quantify the advantage of one estimator with respect to the other in terms of number of measurements that is required to reach a target precision, which amounts to reaching the same value of the above covariance determinant. The values shown for the ratio of the determinants in Fig. 10 can be converted into the ratio of number of measurements by assuming that both covariance determinants in Eq. (47) scale as $1/\sqrt{N_{\rm shot}}$ even in the non-asymptotic regime. Although there can be corrections to this scaling, this can give us in most cases an idea of the proportion of measurement shots one can save by using the best estimator. With this assumption, we get $N_{\rm shot,F}/N_{\rm shot,B}\approx r_{\rm NM}^{2}$, with $r_{\rm NM}$ the ratio of the determinants of two different approaches. Thus, if for instance $r_{\mathrm{\rm NM}}=1.5$, we obtain that $N_{\rm shot,F}/N_{\rm shot,B}\approx 2.25$. Therefore, we get more than a 2×2$\times$2 × reduction in the number of measurements with the Bayesian approach in comparison to the frequentist one to reach the same precision. Going back to the values of Fig. 10, we see that the Bayesian approach can result in a considerable improvement as one goes deep in the non-Markovian limit. Before closing this subsection, it is worth recalling once more that, in practice, the frequentist estimation will never be performed at the three optimal times, and the advantage of the Bayesian approach can be even bigger. In Sec. II of the Supplementary Material, we present a detailed comparison of these estimators with another one in which the shots are evenly distributed between the measurements after evolution times that cover uniformly the whole time interval $T$. As discussed in Sec. II of the Supplementary Material, the Bayesian approach is preferable for most of the parameter values, and can again show a big advantage as one enters into the non-Markovian regime.
Discussion